# DanceAnyWay: Synthesizing Beat-Guided 3D Dances with Randomized Temporal Contrastive Learning

### v2

@@ Introduction
[…]
@@ Appendix B Perceptual Study Setup
1 Very Unnatural
Broken dance motions, Broken arms or legs, torso at impossible angles, etc.
Poor
No or arbitrary movements
2 Not Realistic
Unnatural movements, freezing of dance motions, self-collisions, mismatched proportions, etc.
Below Average
Most of the movements appear to be out-of-sync from the audio
3 Looks OK
No serious problems, but the motion does not look very appealing
Average
Some movements are in sync, and some movements are out-of-sync
4 Looks good
No problems and the dance motions look natural
Above Average
Most of the movements appear to be in sync with the audio
5 Looks great!
The dance motions look like they could be from a real person
Excellent!
Movements appear to be well-synced with the audio

### v3

@@ Introduction
Compared to the best baseline of EDGE (Tseng, Castellon, and Liu 2023), our FIDk and FIDg scores are about $12\%$ and $7\%$ better respectively, our $\textrm{MD}_{k}$ and $\textrm{MD}_{g}$ scores are about $4\%$ and $1.5\%$ better respectively, and our BAS and PFS improve by about $22\%$ and $50\%$ respectively. We further demonstrate better beat alignment of our generated dances for a test sample in Fig. 4 and visualize snapshots from the generated sequences in Fig. 5.
@@ Appendix B Perceptual Study Setup
1 Very Unnatural Broken dance motions, Broken arms or legs, torso at impossible angles, etc. Poor No or arbitrary movements
2 Not Realistic Unnatural movements, freezing of dance motions, self-collisions, mismatched proportions, etc. Below Average Most of the movements appear to be out-of-sync from the audio
3 Looks OK No serious problems, but the motion does not look very appealing Average Some movements are in sync, and some movements are out-of-sync
4 Looks good No problems and the dance motions look natural Above Average Most of the movements appear to be in sync with the audio
5 Looks great! The dance motions look like they could be from a real person Excellent! Movements appear to be well-synced with the audio